OpenStudy (anonymous):

How many grams of H2 are needed to react with 2.40g of N2?

OpenStudy (abb0t):

Write the complete balanced chemical equation first. H\(_2\) +N\(_2\) \(\rightarrow\)

OpenStudy (anonymous):

N2(g) + 3H2(g) → 2NH3(g)

OpenStudy (abb0t):

Yes, now next step is to convert from grams of N\(_2\) \(\rightarrow\) H\(_2\). Do you know dimensional analysis?

OpenStudy (abb0t):

Plan: \(\sf \color{}{2.40~g~N_2 \rightarrow mol~N_2 \rightarrow mol~H_2 \rightarrow g~of~H_2}\)

OpenStudy (anonymous):

No, I'm not so good in chemistry that's why I made a account. I've never had chemistry before but I have it as a subject during my education.

OpenStudy (abb0t):

Well, that is the plan, I listed above, that you must take. Do you reemmeber from algebra, cancel out the variables? For instance: \(\sf \color{red}{\frac{ab}{c^2} \times \frac{1}{a} = \frac{b}{c^2}}\) since the a's cancel out? That's basically what you're doing. Except this time, with units. You have \(\sf \color{blue}{2.40~g~N_2 \times\frac{mole~N_2}{mass~of~N_2} \times \frac{3~mol~H_2}{1~mol~H_2}}\). and then continue the same process to find mass.